# AI-Powered Urban Transportation Digital Twin: Methods and Applications

### v1

@@ 5.1.1 Engineering the pipeline
5.1.1.1   Datasets for urban computer vision
Due to limited availability and single perspective of real-world datasets, synthetic data generation from realistic 3D simulators has become increasingly popular for urban computer vision problems. The SYNTHIA dataset, for example, provides a diverse set of images from a simulated city with pixel-level semantic annotations, aiding tasks like semantic segmentation and scene understanding [129]. CARLA has been widely used in object detection and segmentation [115, 59, 101]. Similarly, GTA V has been employed to collect game frames for training autonomous agents [128]. Unreal Engine has gained traction as a rendering engine for various vision tasks [4, 55, 29]. MatrixCity utilizes Unreal Engine 5’s City Sample project to benchmark neural rendering models, focusing on pedestrian- and vehicle-free environments [78]. The Boundless project shows YOLOv8x’s enhanced performance on real-world datasets when trained with Unreal Engine 5-synthesized data [151]. However, mixed-perspective datasets that adapt to diverse deployment conditions remain scarce. Advances in open-vocabulary object detection offer a promising avenue for addressing this gap [92, 103, 24].
5.1.1.2   Sensor fusion with time synchronization
@@
5.1.2 Designing edge-cloud architecture and networking
@@
5.1.3 Physics-informed DT development
@@
The mutual interaction between domain knowledge and AI is necessary. Domain knowledge in traffic flow and road safety that has been developed for decades, provides valuable insights into every stage of the DT pipeline. In particular, it helps inform technological advancement and testbed deployment, data collection, model selection and training, resource allocation and optimization, as well as performance metrics selection. For instance, the understanding of how traffic evolves across time and space and how entities interact with one another at a road junction, could guide to what degree of granularity and precision the semantic segmentation should be done in object detection and tracking, how communication and networking resources should be prioritized, and what models and metrics should be used for risk prediction.
5.1.4 Safety guarantee of AI
@@
DTs, heavily relying on hardware, software, and algorithms, are vulnerable to risks posed by probabilistic events. Failures in sensors, communication channels, or algorithms can result in erroneous or missing inputs to downstream prediction or decision-making, ultimately compromising the feedback loop from the DT to the physical. The rise of AI-empowered DTs introduces additional risks, including adversarial attacks, and trustworthiness in real-world applications. How can DTs be designed with provable safety guarantees, accounting for uncertainties, failures, and attacks? To achieve this, several methods have been proposed, including reachability analysis [12], robust optimization [195], and control barrier/Lyapunov functions [13, 75]. These methods are mathematically rigorous to ensure that DTs generate outputs within safe regimes. With the emergence of AI, adapting these methods to ensure the safety guarantees of AI models is becoming increasingly important. For instance, reachability analysis has been employed in safe reinforcement learning [136, 159]. Robust optimization techniques have also been applied to enhance model robustness towards adversarial attacks [19, 40]. Additionally, control barrier functions and Lyapunov functions have been integrated into neural networks to enforce safety constraints and maintain system stability [44, 22].
5.1.5 Sim-to-real gap closing
What distinguishes a traffic DT from a conventional traffic simulator? Capabilities of generating simulations and predictions consistent with the real world is the key. There is a growing amount of studies focused on sim-to-real policy learning [196, 15, 62], i.e., how optimal policies learned from a DT can be generalized to real-world implementation. The larger the sim-to-real gap, the less reliable the DT is. To the best of our knowledge, there is little research on how to close the sim-to-real gap leveraging transportation DTs, which we believe is the key to transportation applications.
5.2 Open questions: Prototyping DTs for human intelligence
@@
Many studies define a DT as emerging technologies, namely, object detection and tracking with edge-cloud computing and communication. These technological enablers, however, are actually the “eyes” of a DT, while what really distinguishes a DT from traditional simulators lies in its “brain,” the prediction and decision making modules that are capable of extracting patterns, modeling semantics, and making informed decisions drawing upon what has been seen and perceived.
Ultimately, we should produce a next-generation of DT ideally with human-like intelligence but automation capabilities. A DT with comparable human-like intelligence or artificial artificial general intelligence (AGI) should consist of a hierarchical cognitive structure analogous to a human’s biological system, backed up by emerging hardware, software, AI algorithms, and API interfaces. The architecture of a proposed DT could consist of (see Fig. 6):
@@
    Brain: data storage and processing, reasoning and planning, inference and generalization, powered by AI;
@@
    Neural systems: edge-cloud networking and computing backbone, powered by resource allocators;
@@
    Communication: natural language processing, powered by large language models (LLM) and foundation models.
Figure 6: Schematic diagram of a DT analogous to humans.
A growing number of sensors and explosive amounts of data have opened up tremendous opportunities for researchers to apply AI to create, train, evaluate, and streamline traffic DT pipelines, from object detection, tracking, prediction, simulation, to traffic operation and management. The emerging LLM and foundational models have further demonstrated great potential in DTs. LLMs could generate new data for training data augmentation or privacy concerns [165, 191, 21], enhance interactions between human users and the DT system [188], making personalized recommendations [177, 38] and even automate the creation process of DTs [10].
Despite the promising future of AI-powered DTs, application of AI to DTs could face challenges. As opposed to classical statistical methods, AI algorithms are generally black boxes where their inner workings are not transparent to developers nor users. Thus, interpretability or explainability using Shapley value [100], PIDL [32], symbolic regression [102], or Kolmogorov-Arnold networks [97] can potentially reveal to some degree the rationale underlying the predictions. Unlike humans, AI models do not understand causes and heavily rely on correlations to make predictions. Without knowing causality could render AI methods incapable of generalizing to unseen data. Thus, augmenting AIs with causal reasoning and inference could increase its deductive capabilities [130, 132, 133]. In addition, generative AIs could produce hallucination, which might lead to nonphysical predictions or unrealistic decision making. Introducing physics based domain knowledge could help fine tune these models, enhance inductive biases, and generate more meaningful outputs.
@@
Acknowledgment
We would like to thank Mengxuan Liu for her assistance in generating the 3D models.

### v2

@@ V-A1 Engineering the pipeline
I. Sensor fusion with time synchronization
@@
II. Designing edge-cloud architecture and networking
@@ III. Integrated sensing and communication
V-A2 Emerging DT applications
With the increasing demand for urban passenger and goods delivery, emerging technology like urban sensing, electrification, connectivity and autonomy, and robotics have been gradually transforming urban streetscapes, which pose new challenges to the operation and management of urban infrastructure and public space. DTs are crucial to improve urban safety and mobility in infrastructure planning, service operation and management, and ultimately, policy making.
Here we do not aim to enumerate comprehensive urban transportation applications with DTs, since there exist survey papers [13, 14, 4, 5, 6], which summarize those in operation (e.g., anomaly detection and warning, emergency response), maintenance, and mobility (e.g., transit operation, or driving). Instead, we focus on emerging applications in urban settings, and the potentials the DT holds for them. In Fig. 13, emerging use cases are categorized based on required communication latency (x-axis), spatial resolution (y-axis), and data bandwidth (z-axis), respectively.
Figure 13: Urban T-DT use cases.
Stochasticity arising from travel demands, un-predicted VRU movement, and traffic gridlocks requires traffic operators to design adaptive traffic signal control (ATSC). With a large amount of agents (including VRUs, traffic lights) continuously interacting in a stochastic environment, classical optimization tools in deterministic environments could fail to capture such complex decision-making processes. There is a surge in employing learning methods, including reinforcement learning and federated learning, to optimize traffic signals [169, 170, 171, 172, 173, 174, 175, 176, 177]. With an increasing concern in data governance and privacy, federated learning witnesses a growing trend that allows optimizing centralized control while preserving distributed data privacy [178, 179, 180, 164, 181, 182].
Driven by growing urban populations, low-altitude economy aims to exploit vertical space in urban environments, advanced by drones and eVTOL (electric vertical take-off and landing). It will foster wide applications, ranging from urban air mobility, urban logistics, agriculture, emergency service, to infrastructure inspection, surveillance and security. Critical challenges arise in sensing, modeling, predicting urban aerodynamics, and collision avoidance with high-rise buildings and flying objects. Open questions include market demand estimation, infrastructure planning, operation and design. DTs hold the potential to the adoption of these emerging technologies, for its power in simulating emerging demand patterns, modeling their kinetics and dynamics, routing and charging behaviors. To build such DTs, CPS technologies and AI methodologies would facilitate precise sensing and perception [146], reliable and fast communication and networking [183], seamless multimodal coordination (such as drone-truck-mobile robot delivery) [184], as well as safe and optimal decisions [185].
V-A3 AI models
AI methods have been applied to every stage of the DT pipeline. We will point out key challenges of AI methods in transportation.
Physics-informed AI
@@ V-A3 AI models
When it comes to the development of a T-DT, the mutual interaction between domain knowledge and AI is necessary. Domain knowledge in traffic flow and road safety that has been developed for decades, provides valuable insights into every stage of the DT pipeline. In particular, it helps inform technological advancement and testbed deployment, data collection, model selection and training, resource allocation and optimization, as well as performance metrics selection. For instance, the understanding of how traffic evolves across time and space and how entities interact with one another at a road junction, could guide to what degree of granularity and precision the semantic segmentation should be done in object detection and tracking, how communication and networking resources should be prioritized, and what models and metrics should be used for risk prediction.
Safety guarantee of AI
@@
DTs, heavily relying on hardware, software, and algorithms, are vulnerable to risks posed by probabilistic events. Failures in sensors, communication channels, or algorithms can result in erroneous or missing inputs to downstream prediction or decision-making, ultimately compromising the feedback loop from the DT to the physical. The rise of AI-empowered DTs introduces additional risks, including adversarial attacks [186, 187, 188], ethical concerns [189, 190, 191], and trustworthiness in real-world applications [192, 193, 194]. How can DTs be designed with provable safety guarantees, accounting for uncertainties, failures, and attacks? To achieve this, several methods have been proposed, including reachability analysis [195], robust optimization [196], and control barrier/Lyapunov functions [197, 198]. These methods are mathematically rigorous to ensure that DTs generate outputs within safe regimes. With the emergence of AI, adapting these methods to ensure the safety guarantees of AI models is becoming increasingly important. For instance, reachability analysis has been employed in safe reinforcement learning [199, 200]. Robust optimization techniques have also been applied to enhance model robustness towards adversarial attacks [201, 202]. Additionally, control barrier functions and Lyapunov functions have been integrated into neural networks to enforce safety constraints and maintain system stability [203, 204].
Evaluation and validation
As opposed to traditional traffic management systems heavily involved with humans, an AI-powered CPS-enabled traffic management system demands high degree of autonomy, with increased complexity and scale. Since testing traffic management strategies could be unethical and unsafe, DT thus becomes a crucial tool for the test and verification. There does not exist a unified scheme about what to validate, verify, and test in a DT. Here, we would like to decompose this problem into several layers. First, we need to evaluate whether an DT represents its corresponding physical world correctly. Ideally, the digital is expected to be the twin of the physical, accordingly, “Grieves performance test” [1] is a high-level abstract way to compare the difference in the outputs of both the physical and the DT. This is associated with real2sim gap to be defined and elaborated more in the next section.
Second, we need to verify the system behavior of an DT-enabled CPS, and ensure that it performs as desired [205], such as safety, efficiency, accuracy, and timeliness. Such testing is non-trivial, due to the integration of cyber and physical components, as well as their two-way coupling via communication and networking. There are normally four types of tests [205], namely, conformance testing (i.e., whether a system conforms to an expected behavior), robustness testing (i.e., whether a system is robust against stochasticity in environments), fragility testing (i.e., whether the output of a system is robust against perturbation in inputs), and security testing (i.e., whether a system is not affected by cyber-attacks). DTs play important roles in the above tests, while experimental design is crucial to cover comprehensive scenarios leveraging game theory [206] and in recent years, generative models [207].
The presence of human factors along the DT pipeline could complicates the evaluation of cyber-physical-human systems, because of randomness and unpredictability in human behaviors [208]. Humans are not only participants in traffic (as drivers, pedestrians, cyclists), they are also creators and users of DTs. For DTs that involve the feedback to humans, human-in-the-loop test is widely used [209, 210, 208]. Recent years have witnessed a growing trend of using augmented reality and virtual reality to engage humans as pedestrians, cyclists, or scooters in virtual environments without inducing real-world risks [134, 211]. Hardware-in-the-loop testing, such as vehicle-in-the-loop [212], could help test the performance of some physical components “live,” while allowing the rest to be simulated within a DT. Since there does not exist a unified approach for the DT testing, evaluation methods, metrics, testing platforms should be developed to facilitate standardized assessment and validation of AI models.
Benchmark datasets and methods
To advance the application of AI in DT, we must “stand on the shoulders of giants.” In other words, each team does not simply develop one DT for its own internal use. Instead, we hope that an AI-powered DT could be generalized to diverse tasks, transferred to diverse spatiotemporal settings, and shared for co-development among global researchers. Accordingly, we need to standardize application scenarios, benchmark datasets and methods, and unified test environments, for repetitive training and test [213]. Benchmark datasets and methods necessitate performance comparison of any newly proposed AI methods against the state-of-the-art (SOTA) methods. Thus, the transportation community must push to open source data, codes, simulation, algorithms, and results for replicability. For example, standard test platforms, such as Gym [214], Flower [215], PettingZoo [216, 217], and Ray RLlib [218], are important to benchmark various AI models and methods, which is mostly missing in the transportation community.
Moreover, open mixed-perspective datasets that adapt to diverse deployment conditions remain scarce. Advances in open-vocabulary object detection offer a promising avenue to address this gap [219, 220, 221].
V-B Challenges and open questions
V-B1 Closing real-to-sim-to-real gap
A key challenge in applying models trained in simulated environments to the real-world is domain shift between the training and test environments. Models trained in simulators might likely experience performance degradation when tested in the real world, where the environment includes unpredictable variations that simulators cannot fully replicate. Such a shift – caused by discrepancies between simulated and real-world conditions – can undermine the generalization of models, resulting in reduced performance in scenarios not represented in training. What distinguishes an DT from a conventional computer simulator? We believe the key lies in its capability of generating simulations and predictions consistent with the real world, as well as preserving intervention performance. Accordingly, two gaps exist while establishing a DT, namely, real-to-sim (real2sim) gap (i.e., the deviation of the simulated digital world from the real-world), and sim-to-real (sim2real) gap (i.e., the performance deviation of interventions implemented in the real-world from those simulated in the digital world). The smaller these two gaps are, the closer an DT is to reality.
Although these gaps penetrate through each stage along the DT pipeline, from computer vision, tracking, to prediction, and intervention, sim2real transfer is more studied in object detection [222] and policy learning [223, 224, 225, 226]. The major application area of sim2real transfer in DT is robotic manipulation [227], while that in urban navigation and autonomous driving has witnessed a gradual surge, participially in computer vision [228] and reinforcement learning [229, 230]. Key methods include domain randomization [231, 207], which introduces variations to augment training data and expose the model to a wide range of scenarios; and domain adaptation [232, 233], a technique of finding mappings to transfer data points observed empirically from two different data distributions.
How do we characterize real2sim gap and control such a gap? This boils down to quantifying errors of the digital representation of a physical world. Minimizing the real2sim gap is key to system identification and representation learning. Depending on observability and internal workings, a system could be modeled as white-box, grey-box, or black-box [234]. Domain knowledge could help represent the physical world with higher accuracy. For example, hybrid twin [235, 236] relies on both data-driven and physics-informed AI [237]. Since calibration of a full model could be time-consuming, expensive, and potentially infeasible, the model reduction philosophy has become popular. Digital shadow [11], a digital model with one-way data exchange from the physical to the digital, takes less effort to build, but could fail to update the state of the physical world once feedback is executed. Digital cousin [238] aims not to build a simulation model that replicates the reality exactly, instead, mainly focus on end-to-end gap, namely, from real-world sensing to intervention.
V-B2 Prototyping DTs for human intelligence
@@
Many studies define a DT as emerging technologies, namely, object detection and tracking with edge-cloud computing and communication. These technological enablers, however, are essentially “eyes” of a DT, while what really distinguishes a DT from traditional simulators lies in its “brain,” the prediction and decision making modules that are capable of extracting patterns, modeling semantics, and making informed decisions drawing upon what has been seen and perceived. How do we establish a foundational DT that develops human-like intelligence with machine automation? A DT with comparable human-like intelligence or artificial general intelligence (AGI) should consist of a hierarchical cognitive structure analogous to a human’s neural system, backed up by emerging hardware, software, AI algorithms, and API interfaces. The architecture of a proposed DT could consist of:
@@
    Neural systems: edge-cloud networking and computing backbone, powered by resource allocators [239] and cognitive DT [240];
@@
    Brain: data storage and processing, reasoning and planning, inference and generalization, powered by causal inference and counterfactual analysis [241];
@@
    Communication and reasoning: natural language processing and vision reasoning, powered by generative AI (GenAI) [242].
A T-DT could be deemed as the world model of a transportation manager. A world model is a mental model learned by an AI agent to simulate the evolution of its environment for action planning and reasoning. It is a special type of DT that relies on the agent’s own sensor information. Moreover, it can be embedded into a DT as the AI agent’s internal, abstract representation of the physical world. A world model emerges from the field of robotic learning, and there is an emerging trend to augment a world model with cognitive and reasoning capabilities for more accurate representation and prediction [243]. Such a trend, we strongly believe, must be the pathway for the next-generation of DTs, despite that DTs could be an external representation of a system that would facilitate engineers to monitor, diagnose, control, and manage the system.
In particular, foundational and GenAI models, which are shown to empower cognitive and reasoning architectures of the world model, have demonstrated great potential in DTs [244, 245, 246]. Large language models (LLM) have been used to generate new data for training [247, 248, 249], enhance interactions between human users and the DT system [250], making personalized recommendations [251, 252] and even automate code generation [253] and creation process of DTs [254]. On the other hand, vision language models (VLM) help augment training datasets for robustness, including generation of critical events [255], videos [256, 257] and simulations [258], as well as visual question answering [259, 260]. Thus, we believe that GenAI-powered DT will be the next-generation of DTs for efficient and safe traffic management.
V-B3 Limits in AI
Despite the promising future of AI-powered DTs, application of AI to DTs could face challenges. As opposed to classical statistical methods, AI algorithms are generally black boxes where their inner workings are not transparent to developers nor users. Thus, interpretability or explainability using Shapley value [261], PIDL [262], symbolic regression [263], or Kolmogorov-Arnold networks [264] can potentially reveal to some degree the rationale underlying the predictions. Unlike humans, AI models do not understand causes and heavily rely on correlations to make predictions. Without knowing causality could render AI methods incapable of generalizing to unseen data. Thus, augmenting AIs with causal reasoning and inference could increase its deductive capabilities [265, 266, 267]. In addition, generative AIs could produce hallucination, which might lead to nonphysical predictions or unrealistic decision making. Introducing physics based domain knowledge could help fine tune these models, enhance inductive biases, and generate more meaningful outputs. On the other hand, the emergence of LLM could facilitate the alignment of AI models with human preference. For example, reinforcement learning from human feedback [268] has seen a rapid growth for preferential learning in autonomous driving [269], as well as the generation of more realistic traffic DTs [270].
@@ V-B3 Limits in AI
In a nutshell, transportation applications in urban settings are generally challenging to design, develop, deploy, and test for their potential unsafe and unethical consequences. Thus, AI-empowered DT plays a critical role in effective implementation of these applications. Despite relatively sparse literature in this domain, we review an ensemble body of literature on how to leverage emerging technologies in sensing, communication, edge and cloud computing, for urban traffic management. We hope this paper can serve as a pointer to help researchers and practitioners understand SOTA methods and gaps on the development of DTs; a bridge to initiate conversations across interdisciplinary researchers; and a road map to exploiting potentials of DTs for urban transportation applications.